OpenStudy (anonymous):

What is the slope of a line that is parallel to y = - 2/3 x + 8

OpenStudy (anonymous):

need to know first the slope of -2/3x + 8

OpenStudy (anonymous):

I'm not sure how to find the slope to this one...

OpenStudy (anonymous):

it's already in slope-intercept form; if it's of the form y=mx+b then the slope is m or the number multiplied to x.

OpenStudy (anonymous):

I think the slope is -2/3 :) they are parallels...have the same slope!!!

OpenStudy (anonymous):

So it's -2/3x+b ? I'm kind of new to this so it's taking forever hah

OpenStudy (anonymous):

Thank you jesus

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

The slope=-2/3 :)

OpenStudy (anonymous):

Could you describe the steps for me so I can see how you did it? :)

OpenStudy (anonymous):

yes,one moment...my english is very bad... ;)

OpenStudy (anonymous):

that's completely fine! :)

OpenStudy (anonymous):

The slope is -2/3.Draw a graph using the equation then find the gradient(Slope)

OpenStudy (anonymous):

see this is a rather conceptual question

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

the slope of one line , is the derivative of this line.Do tou know about to derivate? ...Sorry for my english :(

OpenStudy (anonymous):

this slope means that if you going along by the positive axis "x" 2 metres, you will go down 3 metres. If you drawn it ,you can see it better...It,s difficult in the begining...:)

OpenStudy (debbieg):

When the line equation is in y=mx+b form (as your equation above is), then you can read the slope right off the equation. The slope is the value of m, the coefficient of x. E.g.: y=2x+1 has slope m=2 y=-5x+3 has slope m=-5 So that gives you the slope of the line that is parallel. Then the next thing you have to know is that parallel lines have the same slope. E.g.: y=2x+1 is parallel to y=2x+7 and y=2x-9 and anything that looks like y=2x+b for whatever value of b. So when you have a line equation, and you want the slope of a line that is parallel, that just means the SAME slope as the line for which you are given the equation.